# The Self-Organized Criticality Paradigm in Economics & Finance

### v1

@@ 4.3 The ecology of production networks (dynamics)
Still, the above story is all within the realm of economic equilibrium, where profits are maximized, markets clear and all produced goods are consumed. As exogenous shocks buffet the economy, a new equilibrium is assumed to be almost immediately reached, corresponding to the new values of productivities. As already discussed above, this is a heroic assumption that disregards the dynamics of e.g. inventories, stock-outs and bankruptcies. At the very least one should explain the dynamical process through which equilibrium is reached, and how long one should wait before it is reached – if it is ever reached. There is however no consensus on how to model such dynamics.
@@
A simplified, linear approach was proposed by Dessertaine et al. [84] to model the dynamics of small deviations away from equilibrium, where supply-demand and profit imbalances are used as restoring forces towards equilibrium. The resulting equations are of the general form Eq. (3), with $x_{i}=\{p_{i},\pi_{i}\}$ and $N\to 2N$. Interestingly, the smallest negative eigenvalue $-\kappa^{\star}$ of the dynamical matrix $\mathbb{K}$ goes to zero precisely when the smallest positive eigenvalue of $\mathbb{M}$ goes to zero, i.e., when the supply network becomes critical in the sense defined in the previous subsection [84]. As discussed in section 2.1, as the system reaches criticality, convergence towards equilibrium becomes infinitely slow and small exogenous shocks are disproportionately amplified. Hence, as for the Generalized Lotka-Volterra dynamics, the system becomes infinitely fragile as it reaches the critical point, just before the network has to rearrange. In the present context, it means that some firms have to disappear and/or the supply network has to rewire for the economy to remain functional. [^8] On the slow, complex dynamics of rewiring, see e.g. [85].
@@
However, the above “naive”, linear implementation of the out-of-equilibrium dynamics is theoretically inconsistent, as some incontrovertible constraints are violated (for example, consumption cannot exceed amount of available goods). The fully consistent, Agent Based description proposed in [84] goes beyond the scope of the present paper. The results are summarized in the form of the phase diagram shown in Fig. 3. Dessertaine et al. [84] find that, depending on the parameters of the system (strength of the reaction in the face of imbalances, perishability of goods,…) the economy can either (a) completely collapse; or (b) reach the competitive equilibrium after some time, that can become very large; or else (c) enter a state of perpetual disequilibrium, with purely endogenous fluctuations which can be periodic or completely chaotic, corresponding to a coordination breakdown of firms along the supply chain [86].999 The is actually another possibility (d) where some deflationary equilibrium is reached, see [84]. In this last regime, economic equilibrium still exists but is linearly unstable, and is dynamically out-of-reach. This regime is interesting because it offers yet another explanation for “small shocks, large business cycles”: the economy might be permanently in a turbulent state, far from equilibrium, without necessarily being close to a critical point.
@@
Figure 3: Phase Diagram of the Agent Based firm network model of Dessertaine et al. [84]. $x-$axis: strength of forces counteracting supply/demand and profit imbalances; $y-$axis: goods perishability. Leftmost region (a, violet): the economy collapses; middle region (b, blue): the economy reaches equilibrium relatively quickly; righr region (c, yellow): the economy is in perpetual disequilibrium, with purely endogenous fluctuations. The green vertical sliver (d) corresponds to a deflationary equilibrium, see footnote 9 and [84].
@@
The model of Dessertaine et al. [84] therefore suggests two distinct out-of-equilibrium routes to excess volatility: (a) purely endogenous cycles, resulting from over-reactions and non-linearities, or (b) self-organized criticality, i.e. persistence and amplification of exogenous shocks, governed by the proximity of a boundary in parameter space where the competitive equilibrium becomes unstable. While scenario (a) may appear at first sight to be more generic, the economic forces discussed above could make (b) plausible as well. Specific empirical work is needed to distinguish between these two scenarios, which lead to rather different predictions for the structure and dynamics of fluctuations. The SOC scenario suggests in particular that fluctuations should be correlated over long time scales, corresponding to the propagation of “avalanches” over large portions of the production network.
@@ 4.4 Inflation “avalanches”
In a very interesting recent paper, M. Nirei & J. Scheinkman [87] have suggested that repricing events (i.e. firms increasing their price in the face of inflation) are clustered in time – in other words that inflation occurs in waves, or avalanches. The idea that inflation dynamics is intermittent actually dates back to Rob Engle in 1982, and motivated the introduction of the famous ARCH model [88], [^10] See Eq. (17) below. The 1982 Engle paper has, according to Google scholar, 34,000 citations as of today! before it was generalized and applied to volatility in financial markets [89].
Nirei & Scheinkman develop a rather sophisticated equilibrium model to describe the repricing strategy of profit-optimizing firms. Here we simplify the model but keep its essential ingredients. We assume that the inflation rate index is $I(t)>0$. In a mean-field spirit, we do not model the details of the client-supplier network but assume that firms are only sensitive to the overall inflation index. The real (deflated) log-price $p_{i}$ of firm $i$ then evolves, in the absence of repricing, as
@@
In Ref. [91], special attention was devoted to the case $J>1$, where no stationary state can be reached. Instead, an oscillating “business cycle” dynamics appear, where the feedback mechanism described above leads to the synchronisation of bankruptcies and spikes of unemployment, see also [94] for similar effects. One can expect similar inflation cycles to set in for $J>1$ within the context of the Nirei-Scheinkman model, although the introduction of monetary policy would probably significantly affect the results. Still, the question of why $J$ appears to be empirically so close to unity (i.e. why is the system close to critical) is an open question that certainly requires more thought. In that respect, the analogy between economic activity and the critical brain activity, with its occasional seizures, in quite enticing [11, 12, 13, 14].
@@ 5 Self-Organized Criticality in Finance
Perhaps the most striking observation suggesting some form of market fragility is the fact that most major price jumps do not seem to be related to anything particular happening in the world, at least seen from major news feeds like Bloomberg – see e.g. [27, 102, 28, 29, 103].121212 Note that private information should not make prices jump. Traders who want to benefit from private information should execute their orders in a stealthy manner, as clearly illustrated by Kyle’s model [104]. By making the price jump, they would immediately lose their informational advantage. Of course, some news do make prices jump, but most of the time jumps seem to be a result of endogenous feedback loops. Seemingly innocuous shocks get amplified, resulting in micro-, meso- or even macro-liquidity crises (see section 5.2 below). Hence, markets appear to operate “at the edge of chaos”, which would explain why they tend to overreact to small perturbations. The precise mechanism(s) driving markets close to a critical point are however not known – or rather several such mechanisms have been proposed (as reviewed below), but none of them have been accepted on the basis of clear, falsifiable predictions.
There are actually other, non critical explanations for power-laws and long memory. For example, Gabaix et al. [40] argue that the power-law tail of the return distribution is inherited from the power-law distribution of assets under management. [^13] The fact that the distribution of assets is a power-law could itself be a signature of criticality. However, simple plausible models based on random multiplicative growth can easily explain such a behaviour – see e.g. [105]. Large investment firms trade large quantities and should induce large price moves. This idea, however, does not seem to fly: large order sizes are sliced and diced in such a way to be commensurate to the trading volume (see footnote 12). Hence those large orders do not create large return, but rather long-range correlations in the sign of market orders, see [106, 107, 108]. Furthermore, large price moves seem to be related to sudden liquidity shortages, not to outsized incoming volumes, see [109, 110, 28, 111].
@@ 5.1 ARCH & Hawkes processes
Hawkes processes provide an alternative description for self-exciting feedback loops. A Hawkes process describes Poisson events whose intensity increases with the past activity of the process itself (for a review in the context of financial markets, see [101]). Such an increase is mediated by a kernel $\Phi(\tau)$, as for ARCH processes, but now in continuous time. When $g:=\int{\rm d}\tau\,\Phi(\tau)$ becomes larger than unity, the intensity of the process diverges.
@@ 5.3 Competitive market ecology
Following C. Chiarella [128], investors are split into three broad classes: value investors (arbitraging the difference between market price and their estimate of value), trend followers (buying when the price went up, and selling when it went down) and noise traders (with essentially random strategies). Lux & Marchesi add the possibility for investors to switch between trend following and value depending on the relative pay-off of the two strategies. Quite interestingly, these ingredients appear to be enough to generate fat-tailed returns and long-range volatility correlations, which come from the fact that agents can stick to one kind of strategy for very long times – see [129] for more details on this point. The co-existence of trend followers and value traders is further documented empirically in [130, 131, 132, 133] and refs therein.
@@
The attractive nature of the critical point is then quite clear [135, 129, 136]: either $M<M_{c}$ and new investors are lured in by profit opportunities, such that $M\nearrow\,$; or $M>M_{c}$, current investors then make losses and progressively leave the market, until $M\approx M_{c}$. This generic path to criticality is actually quite enticing and lends credence to the SOC scenario for financial markets, see [136] for a review, and [125, 137, 61, 138, 127, 139] for related ideas. As markets become more complex and unpredictable, they also become more fragile, as discussed in section 3.4.
Note finally that a Generalized Lotka-Volterra description of the financial market ecology was proposed long ago by Doyne Farmer [125], along the line of Eq. (7). As market participants trade and impact prices, other participants either benefit or suffer from those price changes. This could provide another justification of market criticality: investment strategies feed onto one another in mutually beneficial or predator-prey relationships, co-evolving into a marginally stable equilibrium, as in section 3.3.
@@ 5.4 Contagion & financial stability
Elucidating the very mechanisms leading to instabilities, failures and system-wide crises in socio-economic systems is crucial to finding remedies, mitigation measures and proposing adequate regulations. Indeed, as we have argued throughout this paper and in Ref. [144], the quest for “efficiency” and “optimality” and the necessity of resilience, i.e. of robustness against small perturbations, may be mutually incompatible.
@@
At this juncture, however, we have many more narratives suggesting critical behaviour in financial markets and economies as a whole than hard empirical evidence for such an interpretation of the pervasive “excess volatility”, or “small shocks, large business cycle” effects routinely observed in these systems. A renewed data-driven effort, like in [90], is certainly needed to convince the economics community of the relevance of fragility to understand excess fluctuations, without having to invoke major exogenous shocks (that are often difficult to identify and substantiate).
@@ Acknowledgements
I wish to thank Doyne Farmer for asking me to write this piece and allowing me to put my ideas together on this exciting topic. My understanding of the subject owes a lot to many conversations over the years with F. Aguirre-Lopez, C. Aubrun, E. Bacry, M. Benzaquen, G. Biroli, G. Bunin, F. Caccioli, D. Challet, C. Colon, Th. Dessertaine, A. Fosset, X. Gabaix, J. Garnier-Brun, J. Gatheral, I. Giardina, S. Gualdi, S. Hardiman, K. Kanazawa, A. Kirman, P. Le Doussal, A. Majewski, R. Marcaccioli, M. Marsili, I. Mastromatteo, M. Mézard, J. Moran, R. Morel, J.-F. Muzy, D. Panja, N. Patil, F. Patzelt, F. Pijpers, M. Potters, M. Rosenbaum, J. Scheinkman, J. Sethna, M. Smerlak, D. Sornette, N. Taleb, M. Tarzia, U. Weitzel, M. Wyart & F. Zamponi.

### v2

@@ 4.3 The ecology of production networks (dynamics)
Still, the above story is all within the realm of economic equilibrium, where profits are maximized, markets clear and all produced goods are consumed. As exogenous shocks buffet the economy, a new equilibrium is assumed to be almost immediately reached, corresponding to the new values of productivities. As already discussed above, this is a heroic assumption that disregards the dynamics of e.g. inventories, stock-outs and bankruptcies. At the very least one should explain the dynamical process through which equilibrium is reached, and how long one should wait before it is reached – if it is ever reached (on this point, see the insightful introductory chapter of Ref. [87]). [^8] Although it lays out a research program which is in my view of fundamental importance, Franklin Fisher’s 1982 book “Disequilibrium foundations of equilibrium economics” [87] only has 882 citations to date. This shows that mainstream economics is still not really interested in these topics, as testified by “state-of-the-art” macroeconomics models which completely disregard out-of-equilibrium effects, see e.g. [88, 89, 90]. For a related discussion, see [91].
@@
There is however no consensus on how to model such dynamics. A simplified, linear approach was proposed by Dessertaine et al. [92] to model the dynamics of small deviations away from equilibrium, where supply-demand and profit imbalances are used as restoring forces towards equilibrium. The resulting equations are of the general form Eq. (3), with $x_{i}=\{p_{i},\pi_{i}\}$ and $N\to 2N$ and a $2N\times 2N$ stability matrix $\mathbb{K}$. Interestingly, the smallest negative eigenvalue $-\kappa^{\star}$ of $\mathbb{K}$ goes to zero precisely when the smallest positive eigenvalue of $\mathbb{M}$ goes to zero, i.e., when the supply network becomes critical in the sense defined in the previous subsection [92]. As discussed in section 2.1, as the system reaches criticality, convergence towards equilibrium becomes infinitely slow and small exogenous shocks are disproportionately amplified. Hence, as for the Generalized Lotka-Volterra dynamics, the system becomes infinitely fragile as it reaches the critical point, just before the network has to rearrange. In the present context, it means that some firms have to disappear and/or the supply network has to rewire for the economy to remain functional. [^9] On the slow, complex dynamics of rewiring, see e.g. [93].
@@
However, the above “naive”, linear implementation of the out-of-equilibrium dynamics is theoretically inconsistent, as some incontrovertible constraints are violated (for example, consumption cannot exceed amount of available goods). The fully consistent, Agent Based description proposed in [92] goes beyond the scope of the present paper. The results are summarized in the form of the phase diagram shown in Fig. 3. Dessertaine et al. [92] find that, depending on the parameters of the system (strength of the reaction in the face of imbalances, perishability of goods,…) the economy can either (a) completely collapse; or (b) reach the competitive equilibrium after some time, that can become very large; or else (c) enter a state of perpetual disequilibrium, with purely endogenous fluctuations which can be periodic or completely chaotic, corresponding to a coordination breakdown of firms along the supply chain [94].101010 The is actually another possibility (d) where some deflationary equilibrium is reached, see [92]. In this last regime, economic equilibrium still exists but is linearly unstable, and is dynamically out-of-reach. This regime is interesting because it offers yet another explanation for “small shocks, large business cycles”: the economy might be permanently in a turbulent state, far from equilibrium, without necessarily being close to a critical point.
@@
Figure 3: Phase Diagram of the Agent Based firm network model of Dessertaine et al. [92]. $x-$axis: strength of forces counteracting supply/demand and profit imbalances; $y-$axis: goods perishability. Leftmost region (a, violet): the economy collapses; middle region (b, blue): the economy reaches equilibrium relatively quickly; righr region (c, yellow): the economy is in perpetual disequilibrium, with purely endogenous fluctuations. The green vertical sliver (d) corresponds to a deflationary equilibrium, see footnote 10 and [92].
@@
The model of Dessertaine et al. [92] therefore suggests two distinct out-of-equilibrium routes to excess volatility: (i) purely endogenous cycles, resulting from over-reactions and non-linearities, or (ii) self-organized criticality, i.e. persistence and amplification of exogenous shocks, governed by the proximity of a boundary in parameter space where the competitive equilibrium becomes unstable. While scenario (i) may appear at first sight to be more generic, the economic forces discussed above could make (ii) plausible as well. Specific empirical work is needed to distinguish between these two scenarios, which lead to rather different predictions for the structure and dynamics of fluctuations. The SOC scenario suggests in particular that fluctuations should be correlated over long time scales, corresponding to the propagation of “avalanches” over large portions of the production network.
@@ 4.4 Inflation “avalanches”
In a very interesting recent paper, M. Nirei & J. Scheinkman [95] have suggested that repricing events (i.e. firms increasing their price in the face of inflation) are clustered in time – in other words that inflation occurs in waves, or avalanches. The idea that inflation dynamics is intermittent actually dates back to Rob Engle in 1982, and motivated the introduction of the famous ARCH model [96], [^11] See Eq. (17) below. The 1982 Engle paper has, according to Google scholar, 34,000 citations as of today! before it was generalized and applied to volatility in financial markets [97], see section 5.1.
Nirei & Scheinkman develop a rather sophisticated equilibrium model to describe the repricing strategy of profit-optimizing firms. Here we simplify the model but keep its essential ingredients. We assume that the inflation rate index is $I(t)>0$. In a mean-field spirit, we do not model the details of the client-supplier network but assume that firms are only sensitive to the overall inflation index. The real (i.e. deflated) log-price $p_{i}$ of firm $i$ then evolves, in the absence of repricing, as
@@
In Ref. [99], special attention was devoted to the case $J>1$, where no stationary state can be reached. Instead, an oscillating “business cycle” dynamics appear, where the feedback mechanism described above leads to the synchronisation of bankruptcies and spikes of unemployment, see also [102] for similar effects. One can expect similar inflation cycles to set in for $J>1$ within the context of the Nirei-Scheinkman model, although the introduction of monetary policy would probably significantly affect the results. Still, the question of why $J$ appears to be empirically so close to unity (i.e. why is the system close to critical) is an open question that certainly requires more thought. In that respect, the analogy between economic activity and the critical brain activity, with its occasional seizures [11, 12, 13, 14], is quite enticing – after all, economic agents are, like neurons, processing information, computing prices and productions, and sending signals to other agents.
@@ 5 Self-Organized Criticality in Finance
Perhaps the most striking observation suggesting some form of market fragility is the fact that most major price jumps do not seem to be related to anything particular happening in the world, at least seen from major news feeds like Bloomberg – see e.g. [27, 110, 28, 29, 111].131313 Note that private information should not make prices jump. Traders who want to benefit from private information should execute their orders in a stealthy manner, as clearly illustrated by Kyle’s model [112]. By making the price jump, they would immediately lose their informational advantage. Of course, some news does make prices jump, but most of the time jumps seem to be a result of endogenous feedback loops. Seemingly innocuous shocks get amplified, resulting in micro-, meso- or even macro-liquidity crises (see section 5.2 below). Hence, markets appear to operate “at the edge of chaos”, which would explain why they tend to overreact to small perturbations. The precise mechanism(s) driving markets close to a critical point are however not known – or rather several such mechanisms have been proposed (as reviewed below), but none of them have been accepted on the basis of clear, falsifiable predictions.
There are actually other, non-critical explanations for power-laws and long memory. For example, Gabaix et al. [40] argue that the power-law tail of the return distribution is inherited from the power-law distribution of assets under management. [^14] The fact that the distribution of assets is a power-law could itself be a signature of criticality. However, simple plausible models based on random multiplicative growth can easily explain such a behaviour – see the discussion in [113] and [42]. Large investment firms trade large quantities and should induce large price moves. This idea, however, does not seem to fly: large order sizes are sliced and diced in such a way to be commensurate to the trading volume (see footnote 13). Hence those large orders do not create large return, but rather long-range correlations in the sign of market orders, see [114, 115, 116, 117]. Furthermore, large price moves seem to be related to sudden liquidity shortages, not to outsized incoming volumes, see [118, 119, 28, 120].
@@ 5.1 ARCH & Hawkes processes
Hawkes processes provide an alternative description for self-exciting feedback loops. A Hawkes process describes Poisson events whose intensity increases with the past activity of the process itself (for a review in the context of financial markets, see [109]). Such an increase is mediated by a kernel $\Phi(\tau)$, as for ARCH processes, but now in continuous time. When $g:=\int{\rm d}\tau\,\Phi(\tau)$ becomes larger than unity, the intensity of the process diverges. When $g$ is precisely equal to $1$ the process is critical and only exists for long memory (power-law) kernels [124].
@@ 5.3 Competitive market ecology
Following C. Chiarella [138], investors are split into three broad classes: value investors (arbitraging the difference between market price and their estimate of value), trend followers (buying when the price went up, and selling when it went down) and noise traders (with essentially random strategies). Lux & Marchesi add the possibility for investors to switch between trend following and value depending on the relative pay-off of the two strategies (see also [139, 140]). Quite interestingly, these ingredients appear to be enough to generate fat-tailed returns and long-range volatility correlations, which come from the fact that agents can stick to one kind of strategy for very long times – see [141] for more details on this point. The co-existence of trend followers and value traders is further documented empirically in [142, 143, 144, 145] and refs therein.
@@
The attractive nature of the critical point is then quite clear [147, 141, 148] and somewhat akin to the classic Grossman-Stiglitz paradox [149]: either $M<M_{c}$ and new investors are lured in by profit opportunities, such that $M\nearrow\,$; or $M>M_{c}$, current investors then make losses and progressively leave the market, until $M\approx M_{c}$. As markets become more complex and unpredictable, they also become more fragile, as discussed in section 3.4.
This generic path to criticality is actually quite enticing and lends credence to the SOC scenario for financial markets, see [148] for a review, and [135, 150, 151, 64, 152] for related ideas, in particular the difficulty for prices to reach equilibrium when incentives provided by mispricing tend to zero [149, 153, 137].
Note finally that a Generalized Lotka-Volterra description of the financial market ecology was derived long ago by Doyne Farmer [135], along the lines of Eq. (7). As market participants trade and impact prices, other participants either benefit or suffer from those price changes. This could provide another justification of market criticality: investment strategies feed onto one another in mutually beneficial or predator-prey relationships, co-evolving into a marginally stable equilibrium, as in section 3.3.
@@ 5.4 Contagion & financial stability
Elucidating the very mechanisms leading to instabilities, failures and system-wide crises in socio-economic systems is crucial to finding remedies, mitigation measures and proposing adequate regulations. Indeed, as we have argued throughout this paper and in Ref. [158], the quest for “efficiency” and “optimality” and the necessity of resilience, i.e. of robustness against small perturbations, may be mutually incompatible, a point forcefully made in [43].
@@
Several plausible mechanisms could thus bring sufficiently often financial markets and economies as a whole close to, or even through a critical point where the system loses stability. At this juncture, however, we have many more narratives than hard empirical evidence for such an interpretation of the pervasive “excess volatility”, or “small shocks, large business cycle” effects routinely observed in these systems. A renewed data-driven effort, like in [98], is certainly needed to convince the economics community of the relevance of fragility to understand excess fluctuations, without having to invoke major exogenous shocks (that are often difficult to identify and substantiate).
@@ Acknowledgements
I wish to thank Doyne Farmer for asking me to write this piece and allowing me to put my ideas together on this exciting topic. His comments on my initial draft have been very useful. My understanding of the subject owes a lot to many conversations with him over the years, and also with many friends and collaborators: F. Aguirre-Lopez, C. Aubrun, E. Bacry, M. Benzaquen, G. Biroli, J. Bonart, G. Bunin, F. Caccioli, D. Challet, C. Colon, Th. Dessertaine, J. Donier, A. Fosset, X. Gabaix, J. Garnier-Brun, J. Gatheral, I. Giardina, M. Gould, S. Gualdi, S. Hardiman, K. Kanazawa, A. Kirman, P. Le Doussal, A. Majewski, R. Marcaccioli, M. Marsili, I. Mastromatteo, M. Mézard, J. Moran, R. Morel, J.-F. Muzy, D. Panja, N. Patil, F. Patzelt, F. Pijpers, M. Potters, M. Rosenbaum, J. Scheinkman, A. Secchi, J. Sethna, M. Smerlak, D. Sornette, N. Taleb, M. Tarzia, U. Weitzel, M. Wyart & F. Zamponi.